OpenStudy (anonymous):

Question 1.) Find the midpoint of the segment with the given endpoints. A(3, –5) and C(2, 9) A. (1, 5.5) B. (0.5, –7) C. (2.5, 2) D. (8, –7) Any help is better than none! I need help on a few more questions if anyone is willing to help me! I need GOOD correct help, or I will fail!

OpenStudy (anonymous):

A(3, -5) and C(2, 9) are the endpoints.

OpenStudy (anonymous):

x-value = 2+(2-3)/2 = 2.5 y-value = -5+(9--5)/2 = 2 therefore answer is C

OpenStudy (anonymous):

Oh I see... Thank you!

OpenStudy (anonymous):

sure :-)